### v1

@@ 1 Introduction
Over the course of U.S. antitrust litigation history, the resolution of a greater number of cases has hinged on market definition, more so than any other significant aspect (Baker,, 2007). Courts in the United States and Europe require competition authorities to define “markets” before assessing competitive effects, as highlighted in various legal cases and statutes. For example, the FTC-DOJ Merger Guidelines have used both the level and the change in Herfindahl-Hirschman Index (HHI)—which depend on market definition—to determine whether a proposed merger should be presumed illegal. [^3] In their recently released 2023 Merger Guidelines, the FTC and DOJ changed the HHI thresholds for triggering structural presumption.
@@
Unfortunately, the use of market definition entails two critical complications. First, the term market definition bears much ambiguity, and there is little consensus on its definition (Glasner and Sullivan,, 2020). A growing contingent of critics have argued for scrapping the market definition altogether and focusing solely on competitive effects analysis, which need not require arbitrary market delineation. [^4] The hostility to market delineation grew from the early 1930s with the development of the theory of monopolistic competition, as product differentiation makes defining a relevant market “an inevitably artificial line-drawing exercise” (Werden,, 1992; Farrell and Shapiro,, 2010). Louis Kaplow, in a series of papers (Kaplow,, 2010, 2011, 2012, 2013, 2015, 2021, 2022), argues that market definition serves no role except to produce market shares, which themselves serve as poor measures of market power, and that the logic of market delineation is inherently circular. While acknowledging the limitations to market delineation, proponents argue that the practice of market delineation has practical merits on theoretical and empirical grounds (Coate and Simons,, 2012; Werden,, 2013; Nocke and Schutz,, 2018; Glasner and Sullivan,, 2020; Miller and Sheu,, 2021; Nocke and Whinston,, 2022; Nocke and Schutz,, 2023). However, the Supreme Court recently reaffirmed that the practice will not be abandoned soon: In Ohio v. American Express, 585 US (2018), the Court disagreed with the plaintiffs’ argument that market definition was not required because they had demonstrated adverse competitive effects. [^5] The Court disagreed and explained that “courts usually cannot properly apply the rule of reason without an accurate definition of the relevant market,” thereby ruling that the plaintiffs failed to establish a prima facie case. See the Court opinion at https://www.supremecourt.gov/opinions/17pdf/16-1454_5h26.pdf. The case clarified when market definition is needed but not how it ought to be defined.
@@
Second, a single antitrust case may be subject to multiple market definitions. [^6] Glasner and Sullivan, (2020) criticizes the common expectation of courts and practitioners that every case should involve one single relevant market as the single market fallacy. Multiple theories of harm may be relevant. Moreover, the popular Hypothetical Monopolist Test may produce multiple market definitions (Baker,, 2007; Davis and Garcés,, 2009; Salop and Moresi,, 2009). Perhaps most importantly, plaintiffs and defendants face uncertainty over which market definition will prevail in court. Thus, scrutinizing the implications of alternative market delineation is a norm rather than an exception. Despite the aforementioned debates and the pervasiveness of sensitivity analysis, attempts to establish a systematic approach have been lacking. This paper aims to fill this gap.
@@
In this paper, I propose a simple framework for conducting an organized sensitivity analysis with respect to market definition. [^7] Sensitivity analysis studies the relative importance of different input factors on the model output. Early advocates such as Leamer, (1985) have recognized the need for organized sensitivity analysis, which has led to the development of the literature on semiparametric models that weaken parametric assumptions or set-identified models that give up on point identification. Leamer, (1985) argues that “[a] fragile inference is not worth taking seriously…. Conclusions are judged to be sturdy only if the neighborhood of assumptions is wide enough to be credible and the corresponding interval of inferences is narrow enough to be useful.” In antitrust settings, plaintiffs and defendants have strong incentives to employ extensive sensitivity analysis to present the robustness of their results with respect to market definitions and determine where to allocate their investigative resources. The main contributions are twofold. First, I model the sensitivity analysis problem by leveraging the partial order nature of candidate markets and use Hasse diagram—a directed acyclic graph for representing a finite partial order—to visualize the sensitivity of metrics of interest with respect to market definitions. My diagrammatic approach facilitates a compact summary of multi-dimensional measures from the sensitivity analysis. [^8] Mathematical diagrams are used in various branches of science to summarize information. The diagrammatic approach to causal inference has proven extremely useful; see, e.g., Pearl, (1995). I provide a novel application of Hasse diagrams for sensitivity analysis. Second, I measure the importance of each firm (or a group of products/firms) using the Shapley value and the Shapley-Shubik power index. [^9] In its original form, the Shapley value is a solution concept in cooperative game theory that finds a fair distribution of gains by measuring each player’s contribution. Shapley-Shubik power index is the Shapley value defined on simple games in which the outcome of a coalition is binary (0 or 1). I interpret the removal of firms from an antitrust market as the firms “participating in a coalition” to generate the final evaluation metric of interest (e.g., Herfindahl-Hirschman index or gross upward pricing pressure). Shapley value allows the analyst to understand the importance of each firm in determining the final conclusion.
@@
I illustrate the framework’s usefulness with an application to the Albertsons/Safeway merger consummated in 2015. Evaluating a supermarket merger requires drawing boundaries on which grocery chains should be considered sufficiently close competitors to traditional supermarkets. Many store formats serve grocery products to consumers but operate on heterogeneous business models. I examine the sensitivity of market concentration measures with respect to the inclusion and exclusion of wholesale club, natural/gourmet, and limited assortment stores. I also decompose the contribution of each major chain in triggering the structural presumption defined by the 2023 FTC-DOJ Merger Guidelines.
@@
This paper contributes to the large literature on antitrust market definitions; see Werden, (1992), Baker, (2007), and Davis and Garcés, (2009) for reviews. While much effort has been dedicated to identifying the “right” market definition, the literature has paid less attention to the sensitivity analysis issue. I fill this gap by developing novel quantitative tools for conducting organized sensitivity analysis with respect to market definition. I find a novel application of Shapley value, a well-known cooperative game theory solution concept primarily applied to study voting power. The use of Shapley value for sensitivity analysis has been increasing in the statistical and machine learning literature (Owen,, 2014; Song et al.,, 2016; Plischke et al.,, 2021; Rozemberczki et al.,, 2022). To the best of my knowledge, quantitative sensitivity analysis with respect to market definitions is new.
@@
My framework also complements empirical tools that hinge on market definitions. While I focus on HHI as an evaluation metric throughout the paper for simplicity, my framework extends to various statistics that may be of interest, including concentration ratio, Herfindahl-Hirschman index (Hirschman,, 1964; Nocke and Schutz,, 2018; Nocke and Whinston,, 2022; Nocke and Schutz,, 2023), merger simulation (Nevo, 2000a, ), compensating marginal cost reduction (Werden,, 1996), upward pricing pressure (Farrell and Shapiro,, 2010; Conlon and Mortimer,, 2021; Miller and Sheu,, 2021; Koh,, 2024). [^10] In principle, upward pricing pressure and compensating marginal cost reductions need not rely on market definition (if reliable diversion ratio data are available). On many occasions, however, they do because diversion ratios need to be estimated. First, especially in the merger screening stage, it is common for economists to specify the demand function as logit so that diversion ratios can be approximated with the proportional-to-share formula, which depends directly on market definition. Second, even flexible demand models (e.g., random-coefficients logit models) can be sensitive to market size parameters or the set of products included in consumers’ consideration set, which relate to market definition (Nevo, 2000b, ). More broadly, the framework may be applied to other settings where the analyst wants to report the sensitivity of the main results with respect to a partially ordered discrete set of assumptions. [^11] For example, a Hasse diagram can be useful in summarizing how the size of the partially identified sets with respect to alternative assumptions on players’ information and solution concept in discrete game models; see, e.g., Koh, (2023). In the field of empirical industrial organization, it is also standard to consider multiple market definitions or explore the implication of different nesting structures in nested logit models. Note that this paper makes no attempts to select the right market definition nor respond to the current debate over market definition.
@@
The rest of the paper is organized as follows. Section 2 presents a model of sensitivity analysis and the use of Hasse diagrams. Section 3 illustrates how the Shapley value can measure the importance of each firm in market definition. Section 4 considers an empirical application. Section 5 concludes.
@@ 2 A Model of Sensitivity Analysis
Expanding the product and/or geographic markets adds more competitors and reduces each firm’s market share. Many popular statistics for evaluating market power and competitive effects of mergers—e.g., concentration ratio, Herfindahl-Hirschman index (both in absolute level and changes from pre-merger level to post-merger level), compensating marginal cost reduction (Werden,, 1996), and upward pricing pressure (Farrell and Shapiro,, 2010)—are functions of market shares, and expanding market definition generally leads to more conservative (lower) estimates. [^12] Again, this statement applies to upward pricing pressure and compensating marginal cost reductions when demand function is specified as logit or CES. For ease of exposition, I focus on a scenario in which the analyst is interested in examining a model’s sensitivity to the exclusion of potential competitors so that expanding the list of firms to be removed from the market increases the evaluation metric.
Let $\mathcal{N}\equiv\{1,2,...,N\}$ be the set of marginal firms under consideration. Each $\omega\in 2^{\mathcal{N}}$ represents an exclusion set defined as the list of firms to be dropped from the market. Function $f:2^{\mathcal{N}}\to\mathbb{R}$, which I refer to the outcome function, returns the value of an evaluation metric. The analyst’s objective is to study the sensitivity of $f$ with respect to changing $\omega$. I provide a running example below.
@@ 2.1 Hasse Diagram Representation of Finite Partial Order
The partial order $(2^{\mathcal{N}},\subseteq)$, where $\subseteq$ represents the inclusion relation, formalizes the idea that a market definition may be broader or narrower than another, but two markets need not be comparable. To graphically represent the partial order, I use a Hasse diagram, a mathematical diagram representing a transitive reduction ordering of a finite partially ordered set.
@@
Figure 1 is a Hasse diagram for $(2^{\mathcal{N}},\subseteq)$ where $\mathcal{N}=\{1,2,3\}$. With nodes defined by the elements of $2^{\mathcal{N}}$ and edges defined by the inclusion relation, the directed acyclic diagram compactly summarizes the hierarchical relationships among the subsets of $\{1,2,3\}$. The bottom node $\emptyset$ represents the broadest market where no firm is excluded. The top node $\{1,2,3\}$ corresponds to the narrowest market in which all firms in $\{1,2,3\}$ are excluded, and only the merging firms $\{A,B\}$ remain. Note that edges are drawn only between nodes that have a covering relationship; for example, there is no direct arrow from $\{1\}$ to $\{1,2,3\}$ since $\{1,2\}$ (or $\{1,3\}$) is positioned between them.
@@ 2.2 Measures of Competitive Effects
It is possible to augment the diagram by layering each node with the value of the outcome function. In a typical application, $f:2^{\mathcal{N}}\to\mathbb{R}$ is monotonic. In Appendix A, I establish the monotonicity of popular statistics (concentration ratios, HHI, upward pricing pressure, and compensating marginal cost reduction) with respect to market definition. [^14] I emphasize again that while upward pricing pressure and compensating marginal cost do not rely on market definition (if reliable diversion ratio data are available), practitioners often rely on logit demand assumptions to approximate diversion ratios using cross-sectional market shares data.
@@
Suppose that there are a total of eight firms with sales $x_{A}=15$, $x_{B}=15$, $x_{C}=10$, $x_{D}=10$, $x_{E}=10$, $x_{1}=9$, $x_{2}=6$, $x_{3}=3$. I assume firms $\{A,B,C,D,E\}$ are always included in the market but test the sensitivity of HHI with respect to the exclusion of Firms 1, 2, and 3. Figure 2 summarizes the sensitivity of pre-merger HHI with $\mathcal{N}=\{1,2,3\}$. Each node reports the level of HHI for a given market definition, and each edge reports the marginal change in HHI as the market definition narrows. For example, the HHI associated with the broadest market definition (bottom node) is 1439. However, excluding Firm 1 increases the HHI by 230 to 1669. Excluding Firms 1, 2, and 3 leads to an HHI of 2083. In general, as shown in Appendix A, excluding more firms in the relevant market increases HHI insofar as the excluded firms’ sales are not too large.
@@
As seen from the example above, Hasse diagrams can organize complex information based on the partial order of market definitions. Although I have assumed a single outcome function $f$, it is possible to augment the diagram with multidimensional information. Yet, the complexity of the diagram quickly rises as the number of firms in $\mathcal{N}$. Specifically, the Hasse diagram for $(2^{\mathcal{N}},\subseteq)$ has $2^{N}$ nodes and $N\cdot 2^{N-1}$ edges. In the following section, I propose using the Shapley value to measure the average contribution of each firm.
@@ 3.1 Shapley Value for Sensitivity Analysis
Shapley value (Shapley,, 1953) is a single-valued solution concept for coalition games that proposes how to distribute the total gains to the players, assuming that they all collaborate. [^15] Shapley value is defined by an axiomatic approach: it is the unique solution concept that satisfies the efficiency, symmetry, null player, and additive properties. See Maschler et al., (2020) for a textbook treatment of the topic. In Definition 2, the function $v(S)$ describes the value of coalition $S\in 2^{\mathcal{N}}$. For each $S\in 2^{\mathcal{N}}$ such that $i\notin S$, $v(S\cup\{i\})-v(S)$ is the marginal surplus that player $i$ brings to the table. The Shapley value of $i$ is the average marginal contribution over all possible $S$ that permits such calculation. Thus, (1) measures the average marginal contribution of each player over the possible different permutations in which the coalition can be formed. A key property of Shapley value is that
@@
i.e., the sum of Shapley values is equal to the total value of the grand coalition. Shapley value measures how much the players can expect from a fair distribution of total surplus generated by the grand coalition.
@@
Recall from Figure 2 that $f(\emptyset)=1439$ and $f(\{1,2,3\})=2083$. Thus, excluding firms $\{1,2,3\}$ from the market increases the level of HHI by $v(\{1,2,3\})\equiv f(\{1,2,3\})-f(\emptyset)=2083-1439=644$. Table 1 reports the Shapley values of each firm. [^16] I use R package CoopGame (Staudacher and Anwander,, 2019) to compute the Shapley values. Note that the Shapley values sum to the value of grand coalition 644. Column “SV Share” says that Firms 1, 2, and 3 account for 44%, 35%, and 21% of the total increment. The share of average contribution measured by Shapley value differs from the sales shares, which are 50%, 33%, and 17% for Firms 1, 2, and 3, respectively. Using the sales shares to explain the importance of each firm in the HHI calculation overstates Firm 1’s contribution and understates Firm 3’s contribution.
@@
Table 1 is significant because it breaks down each firm’s contribution to the final outcome measure of interest. In this example, while sales shares and Shapley values are roughly proportional, using Shapley values offers a distinct advantage. The analyst derives the Shapley values directly from the outcome values, which depend nonlinearly on shares, whereas sales shares merely preserve rankings. In the following, I demonstrate that the utility of Shapley values is even greater when the outcome measure represents a binary decision rule, which is ultimately what matters.

### v2

@@ 1 Introduction
Over the course of U.S. antitrust litigation history, the resolution of a greater number of cases has hinged on market definition, more so than any other significant aspect (Baker,, 2007). Courts in the United States and Europe require competition authorities to define “markets” before assessing competitive effects, as highlighted in various legal cases and statutes. For example, the DOJ-FTC Merger Guidelines have used both the level and the change in Herfindahl-Hirschman Index (HHI)—which depends on market definition—to determine whether a proposed merger should be presumed illegal.
@@
Unfortunately, the use of market definition entails two critical complications. First, the term market definition bears much ambiguity, and there is little consensus on its definition (Glasner and Sullivan,, 2020). A growing contingent of critics has argued for scrapping the market definition altogether and focusing solely on competitive effects analysis, which need not require arbitrary market delineation. [^3] The hostility to market delineation has grown since the 1930s with the rise of monopolistic competition theory, as product differentiation makes defining a relevant market “an inevitably artificial line-drawing exercise” (Werden,, 1992; Farrell and Shapiro,, 2010). Louis Kaplow, in a series of papers (Kaplow,, 2010, 2011, 2012, 2013, 2015, 2021, 2022), contends that market definition is redundant, producing unreliable market shares and relying on circular logic. Despite its limitations, supporters highlight its practical and theoretical merits (Coate and Simons,, 2012; Werden,, 2013; Nocke and Schutz,, 2018; Glasner and Sullivan,, 2020; Miller and Sheu,, 2021; Nocke and Whinston,, 2022; Nocke and Schutz,, 2023). However, the Supreme Court recently reaffirmed that the practice will not be abandoned soon: In Ohio v. American Express, the Court disagreed with the plaintiffs’ argument that market definition was not required because they had demonstrated adverse competitive effects: the Court explained that “courts usually cannot properly apply the rule of reason without an accurate definition of the relevant market,” thereby ruling that the plaintiffs failed to establish a prima facie case (Supreme Court of the United States,, 2018). The case clarified when market definition is needed but not how it ought to be defined.
@@
Second, a single antitrust case may be subject to multiple market definitions. Multiple theories of harm may be relevant (Glasner and Sullivan,, 2020). Moreover, the popular Hypothetical Monopolist Test may produce multiple market definitions (Baker,, 2007; Davis and Garcés,, 2009; Salop and Moresi,, 2009). Perhaps most importantly, plaintiffs and defendants face uncertainty over which market definition will prevail in court. Thus, for both plaintiffs and defendants, scrutinizing the implications of alternative market delineation is a norm rather than an exception. [^4] In the recent FTC v. Kroger case, the FTC proposed two definitions as the relevant product market. First, it defined the “Supermarkets” market, which consists of traditional supermarkets and supercenters (e.g., Walmart and Target). Second, it also defined the “Large-Format Stores” market, which includes not only traditional supermarkets and supercenters but also club stores (e.g., Costco and Sam’s Club), natural food stores (e.g., Whole Foods and Sprouts Farmers Market), and limited assortment stores (e.g., Trader Joe’s and Aldi). The second market definition served to assess the sensitivity of the market concentration analysis to the market definition. The Court accepted both definitions (U.S. District Court, District of Oregon,, 2024). Despite the aforementioned debates and the pervasiveness of sensitivity analysis, attempts to establish a systematic approach have been lacking. This paper aims to fill this gap.
@@
In this paper, I propose a simple framework for conducting a form of organized sensitivity analysis á la Leamer, (1985) with respect to market definition. The main contributions are twofold. First, I model the sensitivity analysis problem by leveraging the partial order nature of candidate markets and use Hasse diagram—a directed acyclic graph for representing a finite partial order—to visualize the sensitivity of metrics of interest with respect to market definitions. My diagrammatic approach facilitates a compact summary of multi-dimensional measures from the sensitivity analysis. Second, I measure the importance of each firm (or a group of products/firms) using the Shapley value and the Shapley-Shubik power index. In its original form, the Shapley value is a solution concept from cooperative game theory that finds a fair allocation of gains among participants in a coalition by measuring each player’s contribution; Shapley-Shubik power index is the Shapley value defined on simple games in which the outcome of a coalition is binary (0 or 1). I interpret the removal of firms from an antitrust market as the firms “participating in a coalition” to generate the final evaluation metric of interest (e.g., Herfindahl-Hirschman index or gross upward pricing pressure). Shapley value allows the analyst to understand the importance of each firm in determining the final conclusion.
@@
I illustrate the framework’s usefulness with an application to the Albertsons/Safeway merger consummated in 2015. Evaluating a supermarket merger requires drawing boundaries on which grocery chains should be considered sufficiently close competitors to traditional supermarkets. Many store formats serve grocery products to consumers but operate on heterogeneous business models. I examine the sensitivity of market concentration measures with respect to the inclusion and exclusion of wholesale club, natural/gourmet, and limited assortment stores. I also decompose the contribution of each major chain in triggering the structural presumption defined by the 2023 DOJ-FTC Merger Guidelines (U.S. Department of Justice and the Federal Trade Commission,, 2023).
@@
This paper contributes to the large literature on antitrust market definitions; see Werden, (1992), Baker, (2007), and Davis and Garcés, (2009) for reviews. While much effort has been dedicated to identifying the “right” market definition, the literature has paid less attention to sensitivity analysis, even though antitrust practitioners routinely rely on it to defend their arguments. I fill this gap by developing novel quantitative tools for conducting organized sensitivity analysis with respect to market definition, which allows practitioners to formalize their sensitivity analysis problems and take a systematic approach. I provide a novel application of Hasse diagrams for sensitivity analysis, which adds to a set of diagrammatic tools economists can use to summarize complex information (e.g., Pearl, (1995)’s causal diagram). Furthermore, I find a novel application of Shapley value, a well-known cooperative game theory solution concept primarily applied to study voting power. The use of Shapley value for sensitivity analysis has been increasing in the statistical and machine learning literature (Owen,, 2014; Song et al.,, 2016; Plischke et al.,, 2021; Rozemberczki et al.,, 2022). To the best of my knowledge, quantitative sensitivity analysis with respect to market definitions is new.
@@
My sensitivity analysis framework also complements empirical tools that hinge on market definitions. While I focus on HHI as an evaluation metric throughout the paper for simplicity (and its popularity), my framework extends to various statistics that may be of interest, including concentration ratio, Herfindahl-Hirschman index (Hirschman,, 1964; Nocke and Schutz,, 2018; Nocke and Whinston,, 2022; Nocke and Schutz,, 2023), merger simulation (Nevo, 2000a, ), compensating marginal cost reduction (Werden,, 1996), upward pricing pressure (Farrell and Shapiro,, 2010; Conlon and Mortimer,, 2021; Miller and Sheu,, 2021; Koh,, 2024). [^5] Upward pricing pressure and compensating marginal cost reductions can bypass market definition with reliable diversion ratio data. However, they often rely on it, as diversion ratios are estimated using market-based formulas (e.g., logit or CES) or demand models sensitive to market size assumptions and consumer consideration sets (Nevo, 2000b, ). More broadly, the framework may be applied to other settings where the analyst wants to report the sensitivity of the main results with respect to a partially ordered discrete set of assumptions.
Note that this paper does not attempt to select the right market definition nor respond to the current debate over market definition. This paper also does not seek to develop a refinement over the current economic tools for analyzing mergers. Instead, I take the current legal standards around market definition as given and propose a framework for sensitivity analysis—which is always carried out—based on classical game-theoretic tools. My framework can help aid robustness analysis and guide resource allocation.
@@
The rest of the paper is organized as follows. Section 2 presents a model of sensitivity analysis and the use of Hasse diagrams. Section 3 illustrates how the Shapley value can measure the importance of each firm in market definition. Section 4 considers an empirical application. Section 5 concludes. The replication package is available at https://doi.org/10.5281/zenodo.16809457.
@@ 2 A Model of Sensitivity Analysis
Expanding the product and/or geographic markets adds more competitors and reduces each firm’s market share. Many popular statistics for evaluating market power and competitive effects of mergers—e.g., concentration ratio, Herfindahl-Hirschman index (both in absolute level and changes from pre-merger level to post-merger level), compensating marginal cost reduction (Werden,, 1996), and upward pricing pressure (Farrell and Shapiro,, 2010)—are functions of market shares, and expanding market definition generally leads to more conservative (lower) estimates. For ease of exposition, I focus on a scenario in which the analyst is interested in examining a model’s sensitivity to the exclusion of potential competitors so that expanding the list of firms to be removed from the market increases the evaluation metric.
Let $\mathcal{N}\equiv\{1,2,...,N\}$ be the set of marginal firms under consideration. Each $\omega\in 2^{\mathcal{N}}$ represents an exclusion set defined as the list of firms to be dropped from the market. Function $f:2^{\mathcal{N}}\to\mathbb{R}$, which I refer to as the outcome function, returns the value of an evaluation metric. The analyst’s objective is to study the sensitivity of $f$ with respect to changing $\omega$. I provide a running example below.
@@ 2.1 Hasse Diagram Representation of Finite Partial Order
The partial order $(2^{\mathcal{N}},\subseteq)$, where $\subseteq$ represents the inclusion relation [^6] An inclusion relation is defined as $A\subseteq B$ if and only if $A$ is a subset of $B$., formalizes the idea that a market definition may be broader or narrower than another, but two markets need not be comparable. To graphically represent the partial order, I use a Hasse diagram, a mathematical diagram representing a transitive reduction ordering of a finite partially ordered set.
@@
Figure 1 is a Hasse diagram for $(2^{\mathcal{N}},\subseteq)$ where $\mathcal{N}=\{1,2,3\}$. With nodes defined by the elements of $2^{\mathcal{N}}$ and edges defined by the inclusion relation, the directed acyclic diagram compactly summarizes the hierarchical relationships among the subsets of $\{1,2,3\}$. The bottom node $\emptyset$ represents the broadest market where no firm is excluded. The top node $\{1,2,3\}$ corresponds to the narrowest market in which all firms in $\{1,2,3\}$ are excluded, and only the merging firms $\{A,B\}$ remain. Note that edges are drawn only between nodes that have a covering relationship; for example, there is no direct arrow from $\{1\}$ to $\{1,2,3\}$ since $\{1,2\}$ (or $\{1,3\}$) is positioned between them. ∎
@@ 2.2 Measures of Competitive Effects
It is possible to augment the diagram by layering each node with the value of the outcome function. In a typical application, $f:2^{\mathcal{N}}\to\mathbb{R}$ is monotonic. In Appendix A, I establish the monotonicity of popular statistics (concentration ratios, HHI, upward pricing pressure, and compensating marginal cost reduction) with respect to market definition. [^9] I emphasize again that while upward pricing pressure and compensating marginal cost do not rely on market definition (if reliable diversion ratio data are available), practitioners often rely on logit (or CES) demand assumptions to approximate diversion ratios using cross-sectional market shares data.
@@
Suppose that there are a total of eight firms with sales $x_{A}=15$, $x_{B}=15$, $x_{C}=10$, $x_{D}=10$, $x_{E}=10$, $x_{1}=9$, $x_{2}=6$, $x_{3}=3$. I assume firms $\{A,B,C,D,E\}$ are always included in the market but test the sensitivity of HHI with respect to the exclusion of Firms 1, 2, and 3. Figure 2 summarizes the sensitivity of pre-merger HHI with $\mathcal{N}=\{1,2,3\}$. [^10] The values are rounded to the nearest integer. Each node reports the level of HHI for a given market definition, and each edge reports the marginal change in HHI as the market definition narrows. For example, the HHI associated with the broadest market definition (bottom node) is 1439. However, excluding Firm 1 increases the HHI by 230 to 1669. Excluding Firms 1, 2, and 3 leads to an HHI of 2083. In general, as shown in Appendix A, excluding more firms in the relevant market increases HHI insofar as the excluded firms’ sales are not too large.
@@
As seen from the example above, Hasse diagrams can organize complex information based on the partial order of market definitions. Although I have assumed a single outcome function $f$, it is possible to augment the diagram with multidimensional information. Yet, the complexity of the diagram quickly rises with the number of firms in $\mathcal{N}$. Specifically, the Hasse diagram for $(2^{\mathcal{N}},\subseteq)$ has $2^{N}$ nodes and $N\cdot 2^{N-1}$ edges. In the following section, I propose using the Shapley value to measure the average contribution of each firm.
@@ 3.1 Shapley Value for Sensitivity Analysis
Shapley value (Shapley, 1953b, ) is a single-valued solution concept for coalition games that proposes how to distribute the total gains to the players, assuming that they all collaborate. [^11] Shapley value is defined by an axiomatic approach: it is the unique solution concept that satisfies the efficiency, symmetry, null player, and additive properties. See Maschler et al., (2020) for a textbook treatment of the topic. It is possible to extend the definition and assign asymmetric weights to players (Shapley, 1953a, ). In Definition 2, the function $v(S)$ describes the value of coalition $S\in 2^{\mathcal{N}}$. For each $S\in 2^{\mathcal{N}}$ such that $i\notin S$, $v(S\cup\{i\})-v(S)$ is the marginal surplus that player $i$ brings to the table. The Shapley value of $i$ is the average marginal contribution over all possible $S$ that permits such calculation. Thus, (1) measures the average marginal contribution of each player over the possible different permutations in which the coalition can be formed. A key property of Shapley value is that
@@
i.e., the sum of Shapley values is equal to the total value of the grand coalition.
@@
Recall from Figure 2 that $f(\emptyset)=1439$ and $f(\{1,2,3\})=2083$. Thus, excluding firms $\{1,2,3\}$ from the market increases the level of HHI by $v(\{1,2,3\})\equiv f(\{1,2,3\})-f(\emptyset)=2083-1439=644$. Table 1 reports the Shapley values of each firm. [^12] I use R package CoopGame (Staudacher and Anwander,, 2019) to compute the Shapley values. To calculate the Shapley value of Firm 1, for instance, note that there are $3!=6$ ways to sequentially form the grand coalition: 1-2-3, 1-3-2, 2-1-3, 3-1-2, 2-3-1, 3-2-1; each sequence represents a feasible path from node $\emptyset$ to node $\{1,2,3\}$ in the Hasse diagram. In the first two cases, Firm 1 is the first to enter the coalition, so the marginal contribution of Firm 1 is $\nu(\{1\})-\nu(\{0\})=229.9751$. In the third and fourth cases, the marginal contribution of Firm 1 is $\nu(\{1,2\})-\nu(\{2\})=291.9501$ and $\nu(\{1,3\})-\nu(\{3\})=263.0744$, respectively. In the last two cases, Firm 1 is the last to enter the coalition, so it is $\nu(\{1,2,3\})-\nu(\{2,3\})=337.9017$. Taking the average over the six permutations gives $281.7693\approx 282$. Note that the Shapley values sum to the value of the grand coalition 644.
@@
  • Notes: Shapley values (SV) are rounded to the nearest integer.
Column “SV Share” says that Firms 1, 2, and 3 account for 44%, 35%, and 21% of the total increment. The share of average contribution measured by Shapley value differs from the sales shares, which are 50%, 33%, and 17% for Firms 1, 2, and 3, respectively. Using the sales shares to explain the importance of each firm in the HHI calculation overstates Firm 1’s contribution and understates Firm 3’s contribution. ∎
Table 1 is significant because it breaks down each firm’s contribution to the final outcome measure of interest. In this example, while the sales shares and Shapley values are roughly proportional, using Shapley values offers a distinct advantage. The analyst derives the Shapley values directly from the outcome values, which depend nonlinearly on shares, whereas sales shares merely preserve rankings. In the following, I demonstrate that the utility of Shapley values is even greater when the outcome measure represents a binary decision rule, which is ultimately what matters.